A bus leaves a station at 1 P.M., traveling west at an average rate of 44 mi/h. One hour later a second bus leaves the same station, traveling east at a rate of 48 mi/h. At what time will the two buses be 274 mi apart?
Answer by checkley71(8403)

44T+48(T-1)=274
44T+48T-48=274
92T=274+48
92T=322
T=322/92
T=3.5 HOUR AFTER THE FIRST BUS LEAVES THEY WILL BE 274 MILES APART.
1 PM + 3.5 HOUR=4:30 PM.